Rewrite 12^(x+1) = 5^(x) as an exact value using logarithms.
Answer by lwsshak3(11628)

Rewrite 12^(x+1) = 5^(x) as an exact value using logarithms.
(x+1)log12=xlog5
xlog12+log12=xlog5
xlog12-xlog5=-log12
x(log12-log5)=-log 12
x=-log12/(log12-log5)